The domain of the function flg where f(x, y) = 2xy and g(x, y) = x² + y²is R.
1
Expert's answer
2020-06-21T16:48:36-0400

"f:R\to R \ that \ is\ 
(x,y)\mapsto 2xy\\
g:R\to R \ that\ is \ (x,y)
\mapsto x^2+y^2\\
1) h=f/g=\frac{f}{g}=\frac{2xy}{x^2+y^2}\\
h:R\setminus (0,0)\to R \setminus(
0,0) \ because \ x^2+y^2 \ is \ not \ equal \, to \,(0,0).\\
The\ domain \ of \ h\ is \ R\setminus(0,0).\\"